Two times a number plus the square of the number equals 35.
Find the number.

Answer by richwmiller(9135)

2n+n^2=35
n^2+2n-35=0
(n-5)(n+7)=0
n=5 and n=-7
2*5+5^2=35
10+25=35
check
-14+49=35
check